# AI Conversational Interviewing: Transforming Surveys with LLMs as Adaptive Interviewers

### v1

@@ Abstract
Traditional methods for eliciting people’s opinions face a trade-off between depth and scale: structured surveys enable large-scale data collection but limit respondents’ ability to express unanticipated thoughts in their own words, while conversational interviews provide deeper insights but are resource-intensive. This study explores the potential of replacing human interviewers with large language models (LLMs) to conduct scalable conversational interviews. Our goal is to assess the performance of AI Conversational Interviewing and to identify opportunities for improvement in a controlled environment. We conducted a small-scale, in-depth study with university students who were randomly assigned to be interviewed by either AI or human interviewers, both employing identical questionnaires on political topics. Various quantitative and qualitative measures assessed interviewer adherence to guidelines, response quality, participant engagement, and overall interview efficacy. The findings indicate the viability of AI Conversational Interviewing in producing quality data comparable to traditional methods, with the added benefit of scalability. Based on our experiences, we present specific recommendations for effective implementation.
AI Conversational Interviewing: Transforming Surveys with LLMs as Adaptive Interviewers
Alexander Wuttke1, Matthias Aßenmacher1, Christopher Klamm2, Max M. Lang3, Quirin Würschinger1, Frauke Kreuter1, 1LMU Munich, 2University of Mannheim , 3University of Oxford , Correspondence: a.wuttke@lmu.de
@@ 1 Introduction
A structured survey is the most popular tool for learning people’s thoughts (Groves, 2009; Kertzer and Renshon, 2022; Stantcheva, 2023). These surveys typically gather individual orientations through self-reports, asking respondents to select from predefined options on fixed questions. This method allows for efficient data collection across large populations, producing structured, tabular data that is straightforward to analyze and compare across respondents (Krosnick, 1999; Groves, 2009). Due to these benefits, structured surveys continue to hold a prominent position in both academic and commercial research. Despite their established utility, structured surveys with predefined response options have significant limitations (Schwarz and Hippler, 1987; Kash, 2013). Their static and impersonal nature often leads to respondent fatigue, which can diminish engagement and, consequently, the quality of responses (Krosnick, 1999; Jeong et al., 2023). More critically, the rigid format of these surveys constrains respondents from fully expressing their thoughts, restricting them from offering responses that researchers may not have anticipated (Chang et al., 2021; Esses and Maio, 2002; Reja et al., 2003; Baburajan et al., 2022; Duck-Mayr and Montgomery, 2023). This limitation hampers the discovery of new phenomena and prevents a comprehensive understanding of the full spectrum of people’s attitudes. An alternative to structured surveys is conversational interviewing, sometimes called in-depth or semi-structured or qualitative interviewing (Adeoye-Olatunde and Olenik, 2021; Kallio et al., 2016; Adams, 2015). This approach involves interviewers engaging with respondents in a more open-ended format, allowing them to freely express their thoughts on topics of interest. The dynamic nature of in-depth interviews helps alleviate respondent fatigue and permits the exploration of opinions that might not fit within predefined response options. However, this approach requires skilled interviewers capable of conducting nuanced conversations, which limits its application to small sample sizes due to the associated costs.
@@
So, survey research inherently faces a trade-off between depth and scale: researchers must choose between conducting in-depth explorations with small groups through personalized interviews or gathering extensive but uniform data via structured surveys. However, recent advances in natural language processing (Dubey et al., 2024; Üstün et al., 2024; Workshop et al., 2023; Costello et al., 2024) present new possibilities for addressing this dilemma. The conversational capabilities of instruction-finetuned large language models (LLMs) (Wei et al., 2022; Ouyang et al., 2022) have made them applicable across various academic and industrial domains. Because LLMs can engage in human-like conversations (Cai et al., 2024; di San Pietro et al., 2023; Palmer and Spirling, 2023), they have the potential to assist or even replace human interviewers in conducting in-depth interviews. By eliminating the costly need for human interviewers, LLMs could enable scalable in-depth conversations, potentially resolving the trade-off between depth and scale.
@@
We contribute to the emerging paradigm of AI Conversational Interviewing by conducting the first close-up investigation of its practical implementation and performance (cf. Figure 1).
  • We provide a new comprehensive assessment pipeline of AI performance in conducting in-depth interviews
@@
  • We perform a detailed comparative analysis of AI-conducted versus human-conducted in-depth interviews, offering insights into their respective strengths and limitations regarding adherence to interview protocols and response quality
  • We pre-registered the study to ensure transparency in the research process and have published all collected data and analysis code
  • We derive data recommendations for future implementations of AI Conversational Interviewing techniques.
@@ 2 Related Work
Advances in AI research have facilitated multiple ongoing commercial and academic projects that use LLM-powered chatbots for in-depth, qualitative, or semi-structured interviews, as they are interchangeably called (Chopra and Haaland, 2023; Weidmann et al., 2024). Although these implementations vary, they collectively highlight the potential of LLMs for conducting in-depth interviews. However, several critical questions regarding the precise implementation of AI interviewing remain unresolved.
Qualitative studies have extensively explored best practices for conducting in-person interviews (Adams, 2015), offering insights that can inform LLM-based interviewing. However, some of these guidelines must be adapted to meet the specific needs of LLM interviewing.
Studies in survey methodology have extensively examined how different interview implementations influence responses. One line of research has focused on interviewer and mode effects (Mittereder et al., 2018; Malhotra and Krosnick, 2007). The presence of an interviewer significantly impacts respondents, often leading to greater engagement but also increasing the likelihood of socially desirable responses (Atkeson et al., 2014; West and Blom, 2016). In this vein, studies on conversational interviewing has shown that a more active and flexible interviewer who engages with questions from respondents can improve data quality (Schober and Conrad, 1997; Davis et al., 2024; Mittereder et al., 2018). [^1] Our method is similar to traditional ”conversational interviewing” in that it enhances flexibility during the interview. However, AI Conversational Interviewing differs by highlighting the flexibility of the respondents rather than the interviewer. Another important factor is the input mode. Responses to open-ended questions vary depending on whether they are submitted via text or speech. Text input typically requires more effort, which can result in shorter but more carefully considered responses (Gavras et al., 2022; Höhne et al., 2024). Consequently, whether AI interviewing uses text or speech input, the responses will not necessarily be better or worse, but they will differ predictably, as text- and speech-based interviews elicit distinct psychological reactions from participants (Gavras et al., 2022).
@@ 3 Study Design and Implementation
Our study pursues two goals:
  1. 1.
    Assess the performance of AI Conversational Interviewing (in comparison to human-led interviewing)
  2. 2.
    Identify problems and opportunities for improvement of AI Conversational Interviewing
@@
We conducted both AI-led and human-led interviews as part of a class activity, where students were randomly assigned to serve as either interviewers or respondents in the respective conditions. Identical questionnaires were used in both interview settings. After the interview respondents filled out a structured questionnaire to evaluate their interview experience. In the AI interview condition, students monitored the interviewees in real-time to identify any technical issues.
@@ 3.2 Model setup
We implemented a voice-assisted AI Conversational Interviewing pipeline with GPT-4222GPT-4 turbo, version: 04/2024 and a Chainlit-based user interface, using the following task-adapted prompts:
@@
    the system instruction to act as an interviewer;
@@
    the user instructions with specific guidelines, derived from the qualitative literature on human in-depth interviewing (Adams, 2015), specifying desirable and undesirable interviewer behavior (cf. Appendix C);
@@
    a task questionnaire on politics and democracy, developed by a democracy researcher among the authors (cf. Appendix D).
@@ 3.5 Evaluation Metrics
During the AI in-depth interview, one student from each pair was assigned to observe the other student’s interaction with the AI interviewer. The observer was given a form to document any technical difficulties or other issues the respondent encountered during the interview (cf. Appendix F).
@@ 4 Findings
We collected data on six human-led and five AI-conducted in-depth interviews. Human-led interviews were audio-recorded and then transcribed.
@@
Qualitative inspection of the transcribed data shows that both the AI and human interviewers faithfully followed the provided questionnaire. Manual coding of all interviewer behavior shows that neither humans nor AI always acted in full accordance with the interview guidelines (Figure 6). Summarizing across all coded categories, we counted 72 violations per AI interview and 64 violations (-11.11%absent-$\downarrow{\text{-11.11\%}}$↓ -11.11%) per human interview, on average.
@@
Turning from the interviewer’s behavior to the participants’ responses, we see that both interviewing settings succeeded in eliciting answers from respondents at substantial lengths. In the AI interviewer setting, the average response length was 52.39 words. In the human interview setting, the average response length was 32.81 words (-62.63%absent-$\downarrow{\text{-62.63\%}}$↓ -62.63%).
While participants‘ answers to the AI interviewer were substantial in length, were they also meaningful in substance? The transcribed responses were given to human coders to rate response quality. While we observe minor differences across setting, overall, the ratings indicate a similar response quality (Figure 3). Responses in human and AI interviews were rated as similarly clear (i.e., easy to understand), empathetic (i.e., sensitive towards the interviewer), engaged (i.e., high level of enthusiasm or interest), complex (i.e., advanced vocabulary), grammatically correct (i.e., error-free), specific (i.e., detailed information), and adequate in tone (i.e., suitable for the context).
@@
Further qualitative observation indeed suggests that text-based inputs encourage respondents to think before writing, whereas audio recording tends to prompt respondents to “think out loud”, allowing them to develop their thoughts while speaking (see Appendix F for an example).
[…]
How did respondents experience the interviews? Participants felt that both the human and AI interviewers were clear in their questions and that each understood their responses (Figure 4). Respondents in both settings left the interview satisfied. However, participants found the AI interview less interesting and were less likely to repeat it, possibly due to the technical problems with the audio recording. While emphasizing that a satisfactory interview hinges on a flawless technical implementation of the interview process, these findings suggest that the absence of a human interviewer does not necessarily need to go along with a deteriorated interview experience for the respondents.
@@ Appendix I Additional Results
AI Interviewer Human Interviewer
[…]
Clarity 4.3 3.9
Empathy 2.6 2.9
Engagement 2.6 3.2
Grammatical correctness 4.3 3.8
Relevance 4.6 4.3
Response complexity (lower is better) 1.9 2.1
Specificity 3.1 3.6
Tone of answers 3.1 3.3
[…]
Tokens per answers 52.39 32.81
Readability (higher is better) 77.66 62.22
[…]
Clarity 1.5 1.9
Interestingness 2.5 3.9
Repeatability 2.5 3.6
Overall Satisfaction 3.8 3.8
Understanding 4.0 4.3
Table 1: Comparison of AI-conducted vs human-conducted interviews: Qualitative assessments, quantitative measurements, and participant survey results.
@@
AI Interview – Respondent <<<—–>>> Human Interview – Interviewer
AI Interview – Coding <<<—–>>>    Human Interview – Respondent
@@
AI Interview – Respondent <<<—–>>> Human Interview – Interviewer
AI Interview – Coding <<<—–>>>    Human Interview – Respondent

### v2

@@ Abstract
Traditional methods for eliciting people’s opinions face a trade-off between depth and scale: structured surveys enable large-scale data collection but limit respondents’ ability to voice their opinions in their own words, while conversational interviews provide deeper insights but are resource-intensive. This study explores the potential of replacing human interviewers with large language models (LLMs) to conduct scalable conversational interviews. Our goal is to assess the performance of AI Conversational Interviewing and to identify opportunities for improvement in a controlled environment. We conducted a small-scale, in-depth study with university students who were randomly assigned to a conversational interview by either AI or human interviewers, both employing identical questionnaires on political topics. Various quantitative and qualitative measures assessed interviewer adherence to guidelines, response quality, participant engagement, and overall interview efficacy. The findings indicate the viability of AI Conversational Interviewing in producing quality data comparable to traditional methods, with the added benefit of scalability. We publish our data and materials for re-use and present specific recommendations for effective implementation.
AI Conversational Interviewing:
Transforming Surveys with LLMs as Adaptive Interviewers
Alexander Wuttke1, Matthias Aßenmacher1,2, Christopher Klamm2, Max M. Lang3, Quirin Würschinger1, Frauke Kreuter1,2 1LMU Munich, 2Munich Center for Machine Learning (MCML), 3University of Mannheim, 4University of Oxford Correspondence: a.wuttke@lmu.de
@@ 1 Introduction
Structured surveys are popular tools to assess public opinion (Groves, 2009; Kertzer and Renshon, 2022; Stantcheva, 2023). These surveys typically gather individual orientations through self-reports, asking respondents to select from predefined options on fixed questions. This method allows for efficient data collection across large populations, producing structured, tabular data that is straightforward to analyze and comparable across respondents (Krosnick, 1999; Groves, 2009). Due to these benefits, structured surveys hold a prominent position in both academic and commercial research.
Despite their established utility, structured surveys with predefined response options have significant limitations (Schwarz and Hippler, 1987; Kash, 2013). Their static and impersonal nature often leads to respondent fatigue, which can diminish engagement and, consequently, the quality of responses (Krosnick, 1999; Jeong et al., 2023). More critically, the rigid format of these surveys constrains respondents from fully expressing their thoughts, restricting them from offering responses that researchers may not have anticipated (Chang et al., 2021; Esses and Maio, 2002; Reja et al., 2003; Baburajan et al., 2022; Duck-Mayr and Montgomery, 2023).
@@
This limitation hampers the discovery of new phenomena and prevents a comprehensive understanding of the full spectrum of people’s attitudes. An alternative to structured surveys is conversational interviewing, sometimes called in-depth or semi-structured or qualitative interviewing (Adeoye-Olatunde and Olenik, 2021; Kallio et al., 2016; Adams, 2015). It involves interviewers engaging with respondents in a more open-ended format, allowing them to freely express their thoughts on topics of interest. The dynamic nature of conversational interviews helps alleviate respondent fatigue and permits the exploration of opinions beyond predefined response options. However, this approach requires skilled interviewers capable of conducting nuanced conversations, which limits its application to small sample sizes due to the associated costs.
So, survey research faces a trade-off between depth and scale: researchers must choose between conducting in-depth explorations with small groups through or large-scale but rigid surveys. However, recent advances in natural language processing (Dubey et al., 2024; Üstün et al., 2024; Workshop et al., 2023; Costello et al., 2024) present new possibilities for addressing this dilemma. The conversational capabilities of instruction-finetuned large language models (Wei et al., 2022; Ouyang et al., 2022) have made them applicable across various academic and industrial domains. Because LLMs can engage in human-like conversations (Cai et al., 2024; di San Pietro et al., 2023; Palmer and Spirling, 2023), they have the potential to assist or even replace human interviewers in conducting conversational interviews. By eliminating the costly need for human interviewers, LLMs could enable scalable in-depth conversations, potentially resolving the trade-off between depth and scale.
@@
We contribute to the emerging paradigm of AI Conversational Interviewing by conducting the first close-up investigation of its practical implementation and performance (cf. Figure 1):
  • We provide a new comprehensive assessment pipeline of AI performance in conducting conversational interviews
@@
  • We are the first to explore the performance of voice-assisted LLM-based interviewing
  • We are the first to perform a detailed comparative analysis of AI-conducted versus human-conducted conversational interviews
  • We pre-registered the study to ensure transparency in the research process
@@ 2 Related Work
Advances in AI research have facilitated multiple ongoing commercial and academic projects that use LLM-powered chatbots for in-depth, qualitative, or semi-structured interviews, as they are interchangeably called (Chopra and Haaland, 2023; Weidmann et al., 2024). Although implementations vary, the studies collectively highlight the potential of LLMs for conducting conversational interviews. Yet, critical questions regarding the implementation remain unresolved and little is known about the relative performance compared with human-led interviews.
Qualitative studies have extensively explored best practices for conducting in-person interviews (Adams, 2015). Our approach is to build on these insights when implementing AI Conversational Interviewing.
Studies in survey methodology have extensively examined how different interview implementations influence responses. One line of research has focused on interviewer and mode effects (Mittereder et al., 2018; Malhotra and Krosnick, 2007). The presence of an interviewer significantly impacts respondents, often leading to greater engagement but also increasing the likelihood of socially desirable responses (Atkeson et al., 2014; West and Blom, 2016). In this vein, studies on conversational interviewing has shown that a more active and flexible interviewer who engages with questions from respondents can improve data quality (Schober and Conrad, 1997; Davis et al., 2024; Mittereder et al., 2018). [^1] Our method is similar to traditional ”conversational interviewing” in that it enhances flexibility during the interview. However, AI Conversational Interviewing differs by highlighting the flexibility of the respondents rather than the interviewer. Another important factor is the input mode. Responses to open-ended questions vary depending on whether they are submitted via text or speech. Text input typically requires more effort, which can result in shorter but more carefully considered responses (Gavras et al., 2022; Höhne et al., 2024). So, the responses will not necessarily be better or worse depending on input mode, but they will differ predictably, as text- and speech-based interviews elicit distinct psychological reactions from participants (Gavras et al., 2022).
@@ 3 Study Design and Implementation
Our study pursues two goals: (a) Assess the performance of AI Conversational Interviewing (in comparison to human-led interviewing) and (b) Identify problems and opportunities for improvement of AI Conversational Interviewing.
@@
We conducted both AI-led and human-led interviews as part of a class activity, where students were randomly assigned to serve as either interviewers or respondents in the respective conditions. Identical questionnaires were used in both interview settings. After the interview, respondents filled out a structured questionnaire to evaluate their interview experience. In the AI interview condition, students monitored the interviewees in real-time to identify any technical issues.
[…]
@@ 3.2 Model setup
We implemented a voice-assisted AI Conversational Interviewing pipeline with GPT-4222GPT-4 turbo, version: 04/2024 and a Chainlit-based user interface, using the following task-adapted prompts (cf. Appendix C):
@@
    The system instruction to act as an interviewer (You are a survey interviewer named ’InterviewGPT’, an AI interviewer, wanting to find out more about people’s views […])
@@
    the user instructions with specific guidelines, derived from the qualitative literature on human in-depth interviewing (Adams, 2015), specifying desirable and undesirable interviewer behavior ([…] Make sure that your questions do not guide or predetermine the respondents’ answers in any way. Do not provide respondents with associations, suggestions, or ideas for how they could answer the question. […])
@@
    a task questionnaire on politics and democracy, developed by a democracy researcher among the authors (e.g. And what do you think “politics” is? How would you define this term?)
@@ 3.5 Evaluation Metrics
During the AI interview, one student from each pair was assigned to observe the other student’s interaction with the AI interviewer. The observer was given a form to document any technical difficulties or other issues the respondent encountered during the interview (cf. Appendix F).
[…]
@@ 4 Findings
We collected data on six human-led and five AI-conducted interviews. Human-led interviews were audio-recorded and then transcribed.
@@
Qualitative inspection of the transcribed data shows that both the AI and human interviewers faithfully followed the provided questionnaire. Manual coding of all interviewer behavior shows that neither humans nor AI always acted in full accordance with the interview guidelines (Figure 6). Summarizing across all coded categories, we counted 72 violations per AI interview and 64 violations (-$\downarrow{\text{-11.11\%}}$%) per human interview, on average.
@@
Turning from the interviewer’s behavior to the participants’ responses, we see that both interviewing settings succeeded in eliciting answers from respondents at substantial lengths. In the AI interviewer setting, the average response length was 52.39 words. In the human interview setting, the average response length was 32.81 words (-$\downarrow{\text{-62.63\%}}$%).
While participants‘ answers to the AI interviewer were substantial in length, were they also meaningful in substance? The transcribed responses were given to human coders to rate response quality. While we observe minor differences across setting, overall, the ratings indicate a similar response quality. Responses in human and AI interviews were rated as similarly clear (i.e., easy to understand), empathetic (i.e., sensitive towards the interviewer), engaged (i.e., high level of enthusiasm or interest), complex (i.e., advanced vocabulary), grammatically correct (i.e., error-free), specific (i.e., detailed information), and adequate in tone (i.e., suitable for the context).
@@
Although unintended, this presents an analytical opportunity to explore differences between written and audio-recorded responses in the AI interviewer setting. As the survey-methodological literature suggests, the answers of respondents who relied on text input were significantly shorter (on average, 21 tokens per answer) than the answers by respondents who used audio-recorded throughout the AI interview (63 tokens per answer ($\uparrow{\text{+67\%}}$%). So, response length markedly varied with input mode.
@@ 6 Conclusion
Ethics Statement
We affirm that our research adheres to the ACL Ethics Policy. To protect participant privacy, we ensure that no individuals are identifiable. To maximize the public value of our work, we make all underlying data and source code openly available for reuse. We declare that no conflicts of interest could influence the study’s outcomes, interpretations, or conclusions. All funding sources supporting this research are acknowledged in the acknowledgments section. Furthermore, we have rigorously documented our methodology, experiments, and results to enhance the replicability of our findings.
Acknowledgements
We thank Laura Kiemes and Valeriya Barakhvostova for excellent research assistance. We are grateful for helpful comments we received at I2SC Saarbrücken Kick-Off Event, the Mainz Workshop on Citizen Perspectives on Demcracy, and the LMU MCMP seminar series. Alexander Wuttke was funded by LMU’s Young Researcher Support Fund. Matthias Aßenmacher was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the National Research Data Infrastructure – NFDI 27/1 - 460037581.
@@ Appendix I Additional Results
AI Interview – Respondent $<$—–$>$ Human Interview – Interviewer
AI Interview – Coding $<$—–$>$    Human Interview – Respondent
@@
AI Interview – Respondent $<$—–$>$ Human Interview – Interviewer
AI Interview – Coding $<$—–$>$    Human Interview – Respondent